# Better Estimation of the KL Divergence Between Language Models

### v2

@@ 2.3 Control Variate Monte Carlo
A general technique to reduce the variance of any estimator is by introducing control variates (§8.2; Ross, 2002). In the context of estimating the KL divergence between two language models, control variate Monte Carlo was popularized by Schulman (2020) and is commonly implemented in many RLHF libraries (Havrilla et al., 2023; Hu et al., 2024; Sheng et al., 2024). Formally, a control variate is any function $g\colon\Sigma^{*}\to\mathbb{R}$ for which $G\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\operatorname*{\mathbb{E}}\left[g(\boldsymbol{Y})\right]$ can be efficiently computed. [^4] One could also consider control variates of the form $g\colon\Sigma^{*}\to\mathbb{R}^{d}$ for $d>1$ (Geffner and Domke, 2018). Armed with a control variate, we define the control variate Monte Carlo as []
$\!\!\!\delta_{\textsc{cv}}=\frac{1}{M}\!\sum_{m=1}^{M}\!f(\boldsymbol{Y}^{(m)})+\alpha\cdot(g(\boldsymbol{Y}^{(m)})-G).\!\!$ (8)
@@
Consider the control variate Monte Carlo estimator $\delta_{\textsc{cv}}$ defined in Eq. 8, and assume that $\operatorname*{\mathbb{E}}[g(\boldsymbol{Y})]<\infty$. Then $\delta_{\textsc{cv}}$ is an unbiased estimator, and its variance is given by
$\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{cv}}}\right]=\frac{\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]+\alpha^{2}\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]+2\alpha\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]}{M}.$ (9)
@@
Assume $\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]<\infty$. It is straightforward to show that $-\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]/\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]$ is the value for $\alpha$ that minimizes the variance the most. If we plug this optimal value for $\alpha$ in Eq. 9, we see
$\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{cv}}}\right]$ $=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]-\frac{1}{M}\frac{\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]^{2}}{\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]}$ (10a)
$=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]\Big{(}1-\frac{\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]^{2}}{\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]}\Big{)}$ (10b)
@@
$=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]\Big{(}1-\underset{}{\mathrm{Corr}}\left[{\normalcolor f,g}\right]^{2}\Big{)},$ (10c)
@@
which directly translates to reducing the variance of the Monte Carlo estimator (compare Eq. 7a and Eq. 10c). The greater the correlation between $f$ and $g$, the more we reduce the variance. Note that the value of $\alpha$ may be estimated from a pilot sample when it cannot be computed analytically. [^5] Note that the control variate method can be straightforwardly extended to support multiple control variates.
2.4 KL Estimation with a Control Variate
@@
A specific control variate for KL estimation was proposed by Schulman (2020), who defined $g(\boldsymbol{Y})=\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}-1$. Substituting this into Eq. 8, the MC estimator for KL divergence with the control variate is given by
$\!\!\delta_{\textsc{cv}}=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}+\alpha\left(\frac{q(\boldsymbol{Y}^{(m)})}{p(\boldsymbol{Y}^{(m)})}-1\right).$ (11)
This proposal has some notable properties. First, $G=\operatorname*{\mathbb{E}}\left[\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}-1\right]=0$, meaning $G$ is known in advance. [^6] Note that for this to hold, we need have $q(\boldsymbol{y})=0$ whenever $p(\boldsymbol{y})=0$, which is different from the support condition we assumed for $\textsc{kl}(p\mid\mid q)<\infty$. Second, as we prove in Prop. 2, the covariance between $g(\boldsymbol{Y})$ and $f(\boldsymbol{Y})$ is non-positive.
Proposition 2.
@@
Given the random variable $\boldsymbol{Y}\sim p$, let $f(\boldsymbol{Y})=\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$ and $g(\boldsymbol{Y})=\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}-1$. We have $\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]=-\textsc{kl}(p\mid\mid q)-\textsc{kl}(q\mid\mid p)$.
@@
See App. C. ∎
For convenience, Schulman (2020) proposes setting $\alpha$ to $1$. [^7] This estimator is related to Bregman divergence, a connection we explore in App. D. Importantly, since $\log$ is a concave function, we have $\log r\leq r-1$, which means that the estimator is always positive. Note that in cases where the optimal value for $\alpha$ is greater than $1$, setting $\alpha=1$ still reduces variance, albeit not optimally. In general, however, there is no guarantee that this choice of $\alpha$ results in an estimator that is at least as good as the MC estimator in terms of variance.
3 Rao–Blackwellized Monte Carlo
In this article, we propose the application of another classical technique to reduce the variance of the Monte Carlo estimation of the KL divergence—Rao–Blackwellization (Gelfand and Smith, 1990; Casella and Robert, 1996). Despite its standing in the statistics literature, a Rao–Blackwellized Monte Carlo estimator has yet to gain traction in the context of RLHF (von Werra et al., 2020; Havrilla et al., 2023; Hu et al., 2024; Sheng et al., 2024). The starting point of Rao–Blackwellization is the following inequality involving the conditional variance
$\mathop{\mathrm{Var}}\left[{\normalcolor\operatorname*{\mathbb{E}}[\delta\mid T]}\right]\leq\mathop{\mathrm{Var}}\left[{\normalcolor\delta}\right],$ (12)
where $\delta$ is an unbiased estimator of $\operatorname*{\mathbb{E}}[f]$ and $T$ is an additional random variable [^8] Note that $T$ is a function of $\{\boldsymbol{Y}^{(m)}\}_{m=1}^{M}$. We have suppressed this function’s arguments in our notation for improved readability. for which we can explicitly compute $\operatorname*{\mathbb{E}}[\delta\mid T]$. This technique is often referred to as Rao–Blackwellization because the inequality is associated with the Rao–Blackwell theorem (Lehmann and Casella, 1998). Note that in the Rao–Blackwell theorem, we get the stronger result that when $T$ is a sufficient statistic, $\mathop{\mathrm{Var}}\left[{\normalcolor\operatorname*{\mathbb{E}}[\delta\mid T]}\right]$ is optimal, i.e., it is the minimal variance unbiased estimator. However, we can perform Rao–Blackwellization even when $T$ is not sufficient and are still guaranteed that the variance is no worse (Robert and Casella, 2004).
We now apply Rao–Blackwellization to construct an unbiased estimator for KL divergence. We begin by lifting the language model to language process (Meister et al., 2023; Du et al., 2024b). A language process is a sequence $\langle\overline{Y}_{n}\rangle_{n=1}^{\infty}$ of $\overline{\Sigma}$-valued random variables where
$\mathbb{P}(\overline{Y}\!_{n}=\overline{y}\mid\overline{\boldsymbol{Y}}_{<n}=\boldsymbol{y})=\vec{p}(\overline{y}\mid\boldsymbol{y}),$ (13)
for $\boldsymbol{y}\in\Sigma^{*}$ and we further define
$\mathbb{P}(\overline{Y}\!_{n}=\overline{y}\mid\overline{\boldsymbol{Y}}_{<n}=\overline{\boldsymbol{y}})=\begin{cases}1&\text{if }\overline{y}=\textsc{eos}\\ 0&\text{otherwise}\end{cases},$ (14)
where $\overline{\boldsymbol{y}}\in\overline{\Sigma}^{*}\setminus\Sigma^{*}$ and $\overline{\boldsymbol{Y}}_{<n}=\overline{Y}\!_{1}\cdots\overline{Y}\!_{n-1}$, i.e., the concatenation of the first $n-1$ random variables in the language process. We will also write $\overline{\boldsymbol{Y}}$ for the $\overline{\Sigma}^{\infty}$-valued random variable. In the language of Markov chains, eos constitutes an absorbing state, i.e., once the language process generates eos, it cannot generate any other symbol. Moreover, because $p$ is distribution over $\Sigma^{*}$, the set of all finite strings, we are guaranteed that $\lim_{n\rightarrow\infty}\mathbb{P}(\overline{Y}_{n}=\textsc{eos})=1$, i.e., eos is generated by the language process almost surely.
Given the above formulation, we introduce the length-conditioned Monte Carlo estimator of the KL divergence as
@@ 3 Rao–Blackwellized Monte Carlo
$\delta_{\textsc{mc}}^{N}$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}(\overline{\boldsymbol{Y}}_{\leq N}^{(m)})}{\vec{q}(\overline{\boldsymbol{Y}}_{\leq N}^{(m)})}$ (15a)
$=\frac{1}{M}\sum_{m=1}^{M}\log\prod_{n=1}^{N}\frac{\vec{p}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}$ (15b)
$=\frac{1}{M}\sum_{m=1}^{M}\sum_{n=1}^{N}\log\frac{\vec{p}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}$ (15c)
@@
$=\sum_{n=1}^{N}\underbrace{\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}{\vec{q}(\overline{Y}_{n}^{(m)}\mid\overline{\boldsymbol{Y}}^{(m)}_{<n})}}_{\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\delta_{\textsc{mc}}^{n}}$ (15d)
$=\sum_{n=1}^{N}\delta_{\textsc{mc}}^{n}.$ (15e)
In the limit as $N\rightarrow\infty$, the length-conditioned Monte Carlo estimator of the KL divergence converges to the Monte Carlo estimator, defined in Eq. 6a.
Lemma 3.
The length-conditioned MC estimator $\delta_{\textsc{mc}}^{N}$ converges to $\delta_{\textsc{mc}}$ as $N$ goes to $\infty$, i.e., $\lim_{N\rightarrow\infty}\delta_{\textsc{mc}}^{N}=\delta_{\textsc{mc}}$.

### v3

@@ 2.3 Control Variate Monte Carlo Estimation
A general approach to reduce estimator variance is through control variates [32, §8.2]. For KL divergence between language models, this technique was popularized by Schulman [33] and is widely used in RLHF libraries [13, 17, 34]. Formally, a control variate is any function $g\colon\Sigma^{*}\to\mathbb{R}$ for which $G\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\operatorname*{\mathbb{E}}\left[g(\boldsymbol{Y})\right]$ can be efficiently computed. [^7] One could also consider control variates of the form $g\colon\Sigma^{*}\to\mathbb{R}^{d}$ for $d>1$ [10]. We define the control variate Monte Carlo estimator as[]
$\mu_{\textsc{cv}}=\frac{1}{M}\!\sum_{m=1}^{M}\!f(\boldsymbol{Y}^{(m)})+\alpha\cdot(g(\boldsymbol{Y}^{(m)})-G).\!\!$ (5)
where $\alpha\in\mathbb{R}$ is a calibration parameter that must be chosen a priori. The proposition below characterizes the variance of $\mu_{\textsc{cv}}$ as a function of $\alpha$, which will tell use how to choose $\alpha$ optimally.
@@
Consider the control variate MC estimator $\mu_{\textsc{cv}}$ defined in Eq. 5, and assume that $\operatorname*{\mathbb{E}}[g(\boldsymbol{Y})]<\infty$. Then $\mu_{\textsc{cv}}$ is an unbiased estimator, and its variance is given by
$\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{cv}}}\right]=\frac{\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]+\alpha^{2}\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]+2\alpha\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]}{M}.$ (6)
@@
Assume $0<\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]<\infty$. It is straightforward to show that $\alpha^{*}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}-\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]/\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]$ is the value that minimizes the variance. If we plug $\alpha^{*}$ in Eq. 6, and simplify, we see that
@@
$\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{cv}}}\right]=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]\Big(1-\underset{}{\mathrm{Corr}}\left[{\normalcolor f,g}\right]^{2}\Big),$ (7)
@@
which directly translates to reducing the variance of the MC estimator. The magnitude of the correlation between $f$ and $g$ determines the degree of variance reduction. Note that the value of $\alpha^{*}$ may be estimated from a pilot sample when it cannot be computed analytically. [^8] Note that the control variate method can be straightforwardly extended to support multiple control variates.
KL Estimation with a Control Variate.
@@
A specific control variate for KL estimation was proposed by Schulman [33], who defined $g(\boldsymbol{Y})=\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}$. Substituting this into Eq. 5, the MC estimator of the KL divergence with this control variate is
$\!\!\mu_{\textsc{cv}}=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}+\alpha\cdot\left(\frac{q(\boldsymbol{Y}^{(m)})}{p(\boldsymbol{Y}^{(m)})}-1\right).$ (8)
Remarks.
This proposal has some notable properties. First, $G=\operatorname*{\mathbb{E}}\left[\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}\right]=1$, meaning $G$ is known in advance. [^9] Note that for this to hold, we need have $q(\boldsymbol{y})=0$ whenever $p(\boldsymbol{y})=0$, which is different from the support condition we assumed for $\textsc{kl}(p\mid\mid q)<\infty$. Second, $\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]=0$ is zero if and only if $p$ is equal to $q$ (Prop. 8), meaning that when the two distributions are not equal, and $\alpha$ is chosen suitably, we are guaranteed to strictly reduce variance. Schulman [33] proposes setting $\alpha=1$; the benefit of this suboptimal choice is that the resulting estimator is always non-negative (Prop. 7). [^10] Note that $\alpha=1$ is the only value of $\alpha$ that guarantees nonnegativity (see proof of Prop. 7). However, setting $\alpha=1$ will increase variance when $\alpha^{*}<\frac{1}{2}$ (Remark 10). Indeed, our experiments (§ 5.1) confirm that $\alpha=1$ is a poor choice in practice—it is better to estimate $\alpha^{*}$ [^11] Prop. 9 provides the exact conditions on $\alpha$ required for a variance reduction. to ensure that the control variate is correctly calibrated.
3 Rao–Blackwellized Monte Carlo
In this article, we propose the application of another classical technique to reduce the variance of the Monte Carlo estimation of the KL divergence—Rao–Blackwellization [RB; 11, 5]. Despite its standing in the statistics literature, a Rao–Blackwellized Monte Carlo estimator has yet to gain traction in the context of RLHF [40, 13, 17, 34].
We define the Rao–Blackwellized Monte Carlo estimator $\mu_{\textsc{rb}}$ as follows:
$\mu_{\textsc{rb}}$ $\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\frac{1}{M}\sum_{m=1}^{M}\sum_{n=1}^{|\boldsymbol{Y}^{(m)}|}\textsc{kl}(\vec{p}(\cdot\mid\boldsymbol{Y}^{(m)}_{<n})\|\vec{q}(\cdot\mid\boldsymbol{Y}^{(m)}_{<n}))$. (9)
The key benefit of this estimator is that we analytically compute the expectation over the $n\textsuperscript{th}$ symbol in each string rather than relying on the single sampled value at that position.
Rao–Blackwellization Background.
The starting point of Rao–Blackwellization is the following inequality involving the conditional variance: $\mathop{\mathrm{Var}}\left[{\normalcolor\operatorname*{\mathbb{E}}[\mu\mid T]}\right]\leq\mathop{\mathrm{Var}}\left[{\normalcolor\mu}\right]$, where $\mu$ is an unbiased estimator of $\operatorname*{\mathbb{E}}[f]$ and $T$ is a statistic [^12] Note that $T$ is a function of $\{\boldsymbol{Y}^{(m)}\}_{m=1}^{M}$. We have suppressed this function’s arguments in our notation for improved readability. for which we can explicitly compute $\operatorname*{\mathbb{E}}[\mu\mid T]$. This technique is often referred to as Rao-–Blackwellization because the inequality is associated with the Rao–Blackwell theorem [21]. [^13] Note that in the Rao–Blackwell theorem, we get the stronger result that when $T$ is a sufficient statistic, $\mathop{\mathrm{Var}}\left[{\normalcolor\operatorname*{\mathbb{E}}[\mu\mid T]}\right]$ is optimal, i.e., it is the minimal-variance, unbiased estimator. However, we can perform Rao–Blackwellization even when $T$ is not sufficient and are still guaranteed that the variance is no worse [31].
Notation.
Before moving forward, we introduce some convenient notation. Let ${\overline{\boldsymbol{Y}}}^{(m)}$ denotes the eos-padded version of $\boldsymbol{Y}^{(m)}$. Additionally, we extend the definition of the prefix probabilities $\vec{p}$ and $\vec{q}$ to strings containing padding symbols, i.e., $\boldsymbol{y}\notin\Sigma^{*}$, with the following additional case: $\vec{p}(y\mid\boldsymbol{y})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\mathbbm{1}\{y=\textsc{eos}\}$, and $\vec{q}(y\mid\boldsymbol{y})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\mathbbm{1}\{y=\textsc{eos}\}$.
Understanding Our Rao–Blackwellized Estimator.
We now present a proof that our Rao–Blackwellized estimator $\mu_{\textsc{rb}}$ is unbiased and indeed does result in a variance reduction. One might wonder why this requires more than a straightforward application of the Rao–Blackwell theorem. The reason is that $\mu_{\textsc{rb}}$ does not arise directly from the standard formulation of Rao–Blackwellization. Instead, we apply Rao–Blackwellization separately to each summand, where each summand corresponds to an estimator over strings of a particular length. In general, Rao–Blackwellizing the summands pointwise does not guarantee a reduction in the variance of their sum, since the summands may be correlated. However, in the specific case of $\mu_{\textsc{rb}}$, we can prove that the overall variance is reduced, despite Rao–Blackwellizing the summands independently, as we state in the following theorem.
Theorem 2.
Suppose the MC estimator $\mu_{\textsc{mc}}$ has finite variance, i.e., $\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{mc}}}\right]<\infty$. Then the following properties regarding $\mu_{\textsc{rb}}$ hold:
  1. $(i)$
    $\operatorname*{\mathbb{E}}[\mu_{\textsc{rb}}]=\textsc{kl}(p\mid\mid q)$    (unbiasedness)
  1. $(ii)$
@@
    $\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{rb}}}\right]\leq\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{mc}}}\right]$    (variance reduction)
@@
See App. C for a detailed proof. However, we provide the proof sketch below for the reader to quickly understand the structure of the argument, which is broken down into three steps.
  1. (1)
    Step-wise Estimation. We begin by Rao–Blackwellizing the step-wise Monte Carlo estimator for any $n>0$,
@@
    $\mu_{\textsc{mc}}^{n}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}$ $=\frac{1}{M}\sum_{m=1}^{M}\log\frac{\vec{p}({\overline{Y}}_{n}^{(m)}\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n})}{\vec{q}({\overline{Y}}_{n}^{(m)}\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n})}$. (10)
    Intuitively, $\mu_{\textsc{mc}}^{n}$ measures the average KL of the $n\textsuperscript{th}$ symbol. The next step is to define $T_{n}({\overline{\boldsymbol{Y}}})={\overline{\boldsymbol{Y}}}_{<n}$ and apply Rao–Blackwellization to each $\mu_{\textsc{mc}}^{n}$ as follows:
@@
    $\mu_{\textsc{rb}}^{n}$ $\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{(1)},\ldots,{\overline{\boldsymbol{Y}}}^{(M)}}\left[\mu_{\textsc{mc}}^{n}\mid T_{n}\right]$ (11a)
    $=\frac{1}{M}\sum_{m=1}^{M}\operatorname*{\mathbb{E}}_{{\overline{\boldsymbol{Y}}}^{(m)}}\left[\log\frac{\vec{p}({\overline{Y}}_{n}^{(m)}\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n})}{\vec{q}({\overline{Y}}_{n}^{(m)}\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n})}\middle|T_{n}({\overline{\boldsymbol{Y}}}^{(m)})\right]$ (11b)
    $=\frac{1}{M}\sum_{m=1}^{M}\textsc{kl}(\vec{p}(\cdot\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n})\|\vec{q}(\cdot\mid{\overline{\boldsymbol{Y}}}^{(m)}_{<n}))$. (11c)
    Now, it is clear from the Rao–Blackwellization theorem that $\mu_{\textsc{rb}}^{n}$ is unbiased and it provides a variance reduction (i.e., $\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{rb}}^{n}}\right]\leq\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{mc}}^{n}}\right]$) for all $n>0$. Inuitively, the source of the variance reduction in the $\mu_{\textsc{rb}}^{n}$ estimator is that we compute the expectation over the $n\textsuperscript{th}$ symbol exactly rather relying on the sampled symbol at that position.
  2. (2)
    Truncated Estimation. Next, we define $\mu_{\textsc{mc}}^{(N)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{n=1}^{N}\mu_{\textsc{mc}}^{n}$ and $\mu_{\textsc{rb}}^{(N)}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{n=1}^{N}\mu_{\textsc{rb}}^{n}$. Intuitively, these estimators target the KL divergence for symbols up to a maximum length of NN. In Lemma 11, we prove that $\operatorname*{\mathbb{E}}[\mu_{\textsc{rb}}^{(N)}]=\operatorname*{\mathbb{E}}[\mu_{\textsc{mc}}^{(N)}]$, and $\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{rb}}^{(N)}}\right]\leq\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{mc}}^{(N)}}\right]$ for all $N>0$ using the law of total expectation and Jensen’s inequality, where the latter is used in a manner analougsly to have it is used in the original Rao–Blackwellization theorem.
  3. (3)
    Complete Estimation. Now, we consider the complete estimation. First, we observe that the limit of the truncated estimators converge to $\mu_{\textsc{mc}}=\lim_{N\rightarrow\infty}\mu_{\textsc{mc}}^{(N)}$ (Lemma 12), and analogously, $\mu_{\textsc{rb}}=\lim_{N\rightarrow\infty}\mu_{\textsc{rb}}^{(N)}$. Thereby, we are able to show that $\mu_{\textsc{rb}}$ is an unbiased estimator of the KL divergence with variance less than or equal to that of $\mu_{\textsc{mc}}$.
Remarks.
Notably, $\mu_{\textsc{rb}}$ is guaranteed to be non-negative, a property desired by some when designing estimators for the KL divergence between two language models, as the KL divergence itself is always non-negative (cf. remarks in § 2.3). In the case of our Rao–Blackwellized estimator, non-nonegativity follows from the fact that each step-wise estimator computes the exact KL divergence between the two next-symbol distributions, conditioned on the sampled context $\boldsymbol{y}_{<n}^{(m)}$. Since all terms in Eq. 11c are non-negative, $\mu_{\textsc{rb}}$ remains non-negative as well.
Complexity Analysis.
Although computing $\mu_{\textsc{rb}}$ might seem more expensive than $\mu_{\textsc{mc}}$, the overall runtime is dominated by the cost of forward passes. Because a forward pass already involves producing the full distribution over $\overline{\Sigma}$ at each position $n$, the additional $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{O}\!\left(MN|\overline{\Sigma}|\right)$} work required by $\mu_{\textsc{rb}}$ is negligible compared to the $M$ forward passes needed for both $\mu_{\textsc{cv}}$ and $\mu_{\textsc{mc}}$.
4 Estimating the Gradient
KL estimation is essential in many applications, especially in fine-tuning large language models. In reinforcement learning from human feedback (RLHF), for example, the objective includes a KL regularization term to balance reward maximization with staying close to a reference model. Since the language model is a differentiable function of parameters $\boldsymbol{\theta}$ optimized via gradient descent, this setup requires computing the gradient of the KL divergence with respect to $\boldsymbol{\theta}$:
$\boldsymbol{G}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\operatorname*{\mathbb{E}}\left[\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y})}{q(\boldsymbol{Y})}\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\boldsymbol{Y})\right].$ (12)
Therefore, the Monte Carlo estimator of this gradient is
$\boldsymbol{\delta}_{\textsc{mc}}=\frac{1}{M}\sum_{m=1}^{M}\log\frac{p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}\nabla_{\boldsymbol{\theta}}\log p_{\boldsymbol{\theta}}(\boldsymbol{Y}^{(m)}).$ (13)
Now, we derive the Rao–Blackwellized Monte Carlo estimator of the gradient. We start with restating Theorem 2.2 in Malagutti et al. [24], which will prove useful.
Theorem 3 (Malagutti et al. [24]; Theorem 2.2).
Let $p$ and $q$ be language models over $\Sigma$. The following equality holds
where we treat $\vec{p}(\cdot\mid\boldsymbol{y})$ and $\vec{q}\left(\cdot\mid\boldsymbol{y}\right)$ as probability distributions over $\overline{\Sigma}^{*}$.
We refer the reader to Malagutti et al. [24] for the proof. Next, to derive the Rao–Blackwellized estimator of the gradient, we take the gradient of the local KL as stated in the following theorem.
Theorem 4.
Let $p_{\boldsymbol{\theta}}$ and $q$ be two language models over $\Sigma$ and $\vec{p}_{\boldsymbol{\theta}}$ the prefix probability function of $p_{\boldsymbol{\theta}}$. Then, the following equality holds
$\nabla_{\boldsymbol{\theta}}\textsc{kl}(p_{\boldsymbol{\theta}}\mid\mid q)=\sum_{\boldsymbol{y}\in\Sigma^{*}}\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\operatorname*{\mathbb{E}}_{Y}\Big[\log\frac{\vec{p}_{\boldsymbol{\theta}}(Y\mid\boldsymbol{y})}{\vec{q}(Y\mid\boldsymbol{y})}\cdot\left(\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(Y\mid\boldsymbol{y})+\nabla_{\boldsymbol{\theta}}\log\vec{p}_{\boldsymbol{\theta}}(\boldsymbol{y})\right)\Big].$ (15)